Calculate and answer the following question on gene frequencies. Please show how you calculated the gene frequencies!

a. Gene frequency. Much of detecting changes in evolution revolve around tracking gene frequency. You need then to know now to calculate gene frequencies and monitor changes in such. Calculate the following gene frequency.

You sample a population of annual flower in 2002 and find the following phenotype frequencies: Red: 100, Pink: 200, White: 100.

b. You return to the population in 2003 and find the following phenotype frequencies: Red: 100, Pink: 180, White: 81. What changes have occurred in the population with respect to gene frequencies? What factors could explain any differences noted?
Answer #1

Answer:

Color is a codominant trait (i.e. the heterozygotes differ phenotypically from both homozygotes), and the pink individuals are heterozygous.

(a) Frequency of dominant allele (p) = \sqrt{100/400} = 1/2

So, frequency of recessive allele, q = 1 - 1/2 = 1/2

Thus, genotype frequencies: p2 = 0.25 (Red), 2pq = 0.5 (Pink), q2 = 0.25 (White)

(b) Frequency of dominant allele (p) = \sqrt{100/361} = 10/19

Frequency of recessive allele (q) = 1- 10/19 = 9/19

Genotype frequencies: p2 = 0.2767, (Red), 2pq = 0.4975 (Pink), q2 = 0.2237 (White)

So, a comparison of the data obtained in 2002 and 2003 indicates variation in gene frequencies. To judge the significance of the change, a chi-square test can be performed.

If the difference is significant, one or more of the assumptions of the Hardy-Weinberg theorem have been violated, including the following:

Selection may be acting to reduce the frequency of one or more genotypes;

Inbreeding or assortative mating may be reducing the frequency of heterozygotes;

Drift may be reducing the frequency of heterozygotes and increasing the frequency of one of the homozygotes;

Immigration from another population may be contributing genotypes in numbers that skew the expected frequency.